OpenStudy (anonymous):

I need help figuring out this problem. I would prefer an example of it so I can solve the problem myself. thanks -18x^2y^4 over 6x^7y^2

OpenStudy (anonymous):

You need to use the law of exponents here. When you have \[\frac{x^b}{x^a}=x^{b-a}\]

OpenStudy (anonymous):

what would this look like in your problem?

OpenStudy (anonymous):

So do I just add the exponents together

OpenStudy (anonymous):

no. you are not multiplying, you are dividing. When you divide you subtract the exponents of like bases.

OpenStudy (anonymous):

Ok. so I would -18x^2 from 6x^7 correct?

OpenStudy (anonymous):

yeah. the coefficients -18 and 6 are easy, you take care of them like any other division -18/6=-3 Then you take care of the exponents of like bases.

OpenStudy (anonymous):

Ok, so it would be -3x^-5y^2 as my answer?

OpenStudy (anonymous):

yep :) and it is important to recognize that this answer can also be written as \[\frac{-3y^2}{x^5}\]

OpenStudy (anonymous):

but I prefer the way you wrote it

OpenStudy (anonymous):

Thank you so much.

OpenStudy (anonymous):

np, dont forget to medal me to close the question :)

OpenStudy (amistre64):

medaling doesnt close a question; the close button closes a question